# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

### v3

@@ B.2 Function vectors of in-context learning
Appendix C Effects of different performance-critical tokens
@@
To achieve this, we ablate the representations of each performance-critical token, similar to Section 5.2.1. In Section 5.2.1, we assume that the label token is needed for ICL to achieve performance results on par with Standard ICL, as suggested by previous work (Wang et al., 2023). However, it is still not known how the other performance-critical tokens affect the performance without . Hence, we divide our experiments by including or excluding the label tokens to further specifically investigate their effectiveness. We present the results on RTE datasets in Table 12 while full results are shown in Appendix R.
Overall, the above experiments show that the performance-critical tokens should be utilized together to provide the best performance and that removing some of them would cause performance degeneration or instability issues. From the results with , it is observed that all the template tokens (i.e., , , and “:”) contribute to the final performance. Removing one of them would cause a performance degradation. From the results without , the performance becomes less predictable, where adding back a template token (e.g., “:”) does not always bring performance improvements. Moreover, in some datasets, models without can still achieve relatively high performance. These results show that representations of other template tokens may also be seen as information anchors whose representations aggregate and serve information to the final prediction of LLMs, broadening the conclusions of Wang et al. (2023) who claim that only answer tokens serve as information anchors.
Appendix D In-context Learning Templates
@@
Appendix E Representation-level Ablation Examples
We provide a one-shot demonstration of all ablation cases for the representation-level ablation experiments, as shown in Table 15. In these demonstrations, representations of specific token types are masked in the attention mechanism, where <m> denotes that all representations of the token are removed from the attention scope from the test example.
@@ Appendix E Representation-level Ablation Examples
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> |
@@
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
<m> <m> <m> <m> <m> <m> the <m> . <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> the <m> <m> <m> <m> <m> <m> <m> <m> , <m> <m> <m> <m> <m> <m> of <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> into <m> <m> the <m> <m> . |
<m> <m> Technology |
@@
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
<m> <m> <m> First class to <m> moon <m> London - British airline magnate Richard Branson announced a plan on Monday for <m> world’s first commercial space flights <m> saying “thousands” <m> fee-paying astronauts could be sent <m> orbit in <m> near future <m> <m> <m> <m> Technology <m> <m> |
@@
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
<m> <m> <m> First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. <m> <m> <m> Technology <m> <m> |
@@
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: First class to <m> moon <m> London - British airline magnate Richard Branson announced a plan on Monday for <m> world’s first commercial space flights <m> saying “thousands” <m> fee-paying astronauts could be sent <m> orbit in <m> near future <m> <m> Answer: Technology <m> <m> |
@@
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: <m> <m> <m> the <m> . <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> the <m> <m> <m> <m> <m> <m> <m> <m> , <m> <m> <m> <m> <m> <m> of <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> into <m> <m> the <m> <m> . |
@@
Appendix F Experiment Results with More Large Language Models
Experimental results using the Llama 2 and Mistral models are shown in Table 16. The trends observed in these experiments are consistent with those involving the Llama models. These results further reinforce the findings of this paper, indicating that template tokens and stopword tokens are the most prone to serving as performance-critical tokens.
Models | Setting | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB |
Llama 2 7B | Zero-shot | 36.0 | 50.2 | 50.4 | 57.2 | 6.4 | 51.6 | 0.0 |
cont | 1.9 | 0.9 | 61.0 | 50.6 | 12.9 | 48.7 | 53.2 | |
stop | 23.4 | 49.0 | 78.1 | 54.4 | 61.6 | 65.3 | 47.9 | |
temp | 31.3 | 81.1 | 82.6 | 55.2 | 65.5 | 63.9 | 55.4 | |
Llama 2 13B | Zero-shot | 53.3 | 56.2 | 90.8 | 49.0 | 7.6 | 70.0 | 46.4 |
cont | 14.1 | 0.5 | 56.0 | 61.4 | 0.0 | 62.6 | 54.6 | |
stop | 9.6 | 47.2 | 76.8 | 65.2 | 65.3 | 66.5 | 56.8 | |
temp | 18.1 | 78.2 | 93.7 | 62.4 | 70.4 | 71.9 | 52.1 | |
Mistral 7B | Zero-shot | 59.5 | 77.8 | 84.4 | 73.0 | 57.8 | 1.8 | 62.1 |
cont | 10.0 | 43.3 | 52.0 | 66.6 | 10.1 | 64.3 | 60.7 | |
stop | 18.4 | 78.9 | 92.5 | 71.6 | 81.4 | 69.9 | 72.9 | |
temp | 19.7 | 81.7 | 95.9 | 63.9 | 83.3 | 77.9 | 72.5 | |
Llama 2 7B | Standard ICL | 70.7 | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 |
cont | 3.8 | 82.4 | 85.5 | 54.3 | 64.2 | 59.6 | 55.7 | |
stop | 2.5 | 84.8 | 88.0 | 51.7 | 65.7 | 65.8 | 53.2 | |
temp | 32.8 | 0.9 | 61.0 | 50.6 | 12.9 | 48.5 | 53.6 | |
Llama 2 13B | Standard ICL | 73.6 | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 |
cont | 1.3 | 79.0 | 94.1 | 62.7 | 72.4 | 72.1 | 53.6 | |
stop | 2.9 | 80.1 | 89.4 | 61.5 | 74.1 | 69.6 | 49.3 | |
temp | 33.5 | 0.5 | 56.0 | 61.4 | 0.0 | 68.2 | 54.3 | |
Mistral 7B | Standard ICL | 80.2 | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 |
cont | 0.6 | 81.8 | 96.2 | 64.4 | 83.4 | 78.9 | 72.9 | |
stop | 0.8 | 81.3 | 97.0 | 66.5 | 80.5 | 75.5 | 75.7 | |
@@ Appendix F Experiment Results with More Large Language Models
Appendix G Experiment Results with Instruction-tuned Large Language Models
We present the representation-level ablation results of large language models after instruction tuning to confirm that our findings remain consistent. Specifically, we use Llama 2 7B Chat and Llama 2 13B Chat in our experiments. As shown in Table 17, the results align with the findings discussed in Section 5.2.1, with only one exception: the TREC dataset. In this dataset, the input data is structured in a question-answering format (e.g., Who/What/When did …). We hypothesize that, during the supervised fine-tuning process, tokens associated with these question formats may also serve as performance-critical tokens, although they are currently categorized as content tokens in our experiments. Overall, these supplemental results prove further evidence to our findings in the main paper.
Models | Setting | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB |
Llama 2 7B Chat | Zero-shot | - | - | - | - | - | - | - |
cont | 27.9 | 0.7 | 52.6 | 52.2 | 7.9 | 24.8 | 28.9 | |
stop | 72.9 | 77.1 | 90.6 | 60.2 | 75.4 | 66.7 | 67.5 | |
temp | 75.6 | 80.1 | 92.6 | 62.6 | 76.6 | 69.5 | 72.1 | |
Llama 2 13B Chat | Zero-shot | - | - | - | - | - | - | - |
cont | 38.4 | 0.0 | 55.7 | 67.3 | 0.5 | 63.5 | 43.2 | |
stop | 67.1 | 78.5 | 87.6 | 66.9 | 71.4 | 67.2 | 31.1 | |
temp | 72.3 | 82.0 | 93.7 | 65.6 | 72.3 | 72.3 | 47.9 | |
Llama 2 7B Chat | Standard ICL | - | - | - | - | - | - | - |
cont | 76.1 | 80.7 | 93.1 | 62.9 | 76.8 | 70.7 | 72.5 | |
stop | 76.4 | 81.7 | 94.4 | 61.9 | 74.9 | 71.0 | 74.3 | |
temp | 31.4 | 0.7 | 52.6 | 52.2 | 7.9 | 24.1 | 51.1 | |
Llama 2 13B Chat | Standard ICL | - | - | - | - | - | - | - |
cont | 74.7 | 83.4 | 93.6 | 66.4 | 74.3 | 74.6 | 56.1 | |
stop | 69.8 | 78.9 | 94.1 | 57.7 | 72.8 | 70.1 | 45.4 | |
@@ Appendix G Experiment Results with Instruction-tuned Large Language Models
Appendix H Stopword Tokens
For the results shown in the main paper, we used the stopword token (Sarica and Luo, 2021) list shown in Table 10. This list only includes the stopword tokens from the task instruction, aiming to minimize their presence. We made this choice under the assumption that task-affecting information should be stored densely in a few tokens. Hence, the number of tokens whose representations affect the final task performance significantly should be small.
Nevertheless, one might be curious about the results if we used a more complete stopword list. In this case, we utilize a more comprehensive stopword token list of NLTK666https://gist.github.com/sebleier/554280 shown in Table 11 and conduct the representation-level ablation once more. The results are presented in Table 18. It can be observed that all the conclusions from Section 5.2.1 are still well established. A few results are different from Table 2 and Table 3 because we masked the representations of the “</s>” token in this set of experiments. We claim that this masking does not impact the main findings of these experiments.
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB |
OpenLlama 3B | Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 |
+ cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | |
+ stop | 38.0 | 85.1 | 31.6 | 54.6 | 58.8 | 55.7 | |
+ temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | |
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | |
- temp | 42.1 | 87.2 | 25.9 | 56.3 | 58.3 | 57.4 | |
- cont | 57.1 | 88.4 | 27.1 | 62.6 | 56.8 | 52.4 | |
- stop | 61.6 | 90.7 | 24.8 | 62.2 | 56.7 | 51.9 | |
Llama 7B | Zero-shot | 25.0 | 29.2 | 41.4 | 0.0 | 54.2 | 3.6 |
+ cont | 32.4 | 57.9 | 42.5 | 12.5 | 55.5 | 46.1 | |
+ stop | 59.9 | 85.9 | 51.7 | 28.9 | 56.0 | 52.7 | |
+ temp | 70.8 | 90.2 | 58.4 | 66.2 | 66.3 | 73.5 | |
Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | |
- temp | 64.7 | 84.1 | 54.0 | 56.7 | 56.1 | 48.2 | |
- cont | 75.4 | 93.8 | 59.8 | 67.5 | 66.8 | 74.8 | |
- stop | 81.4 | 94.2 | 60.5 | 67.9 | 67.6 | 72.1 | |
Llama 13B | Zero-shot | 59.0 | 18.0 | 37.0 | 0.0 | 0.0 | 0.0 |
+ cont | 30.6 | 52.4 | 43.8 | 13.0 | 60.2 | 45.5 | |
+ stop | 72.7 | 78.7 | 49.2 | 27.4 | 58.5 | 27.1 | |
+ temp | 78.5 | 92.3 | 59.0 | 74.2 | 67.4 | 52.3 | |
Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | |
- temp | 71.7 | 80.1 | 56.2 | 8.7 | 56.5 | 29.3 | |
- cont | 79.3 | 93.4 | 60.1 | 74.1 | 68.4 | 47.6 | |
- stop | 79.2 | 94.1 | 59.3 | 73.8 | 68.9 | 44.6 | |
Llama 33B | Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 |
+ cont | 27.8 | 61.7 | 61.9 | 10.8 | 64.2 | 68.1 | |
+ stop | 74.7 | 93.6 | 66.9 | 70.8 | 69.1 | 63.8 | |
+ temp | 80.6 | 95.2 | 63.1 | 71.9 | 78.7 | 84.0 | |
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | |
- temp | 79.5 | 93.8 | 58.5 | 62.8 | 68.0 | 68.0 | |
- cont | 82.7 | 95.9 | 62.9 | 74.1 | 79.6 | 83.1 | |
- stop | 84.4 | 96.1 | 61.8 | 72.8 | 79.4 | 82.1 |
Appendix I Analysis of the Next-line Tokens

### v4

@@ C.2 Function vectors of in-context learning
Appendix D Effects of different performance-critical tokens
@@
To achieve this, we ablate the representations of each performance-critical token, similar to Section 5.2. In Section 5.2, we assume that the label token is needed for ICL to achieve performance results on par with Standard ICL, as suggested by previous work (wang2023label). However, it is still not known how the other performance-critical tokens affect the performance without . Hence, we divide our experiments by including or excluding the label tokens to further specifically investigate their effectiveness. We present the results on RTE datasets in Table 13 while full results are shown in Appendix S.
Overall, the above experiments show that the performance-critical tokens should be utilized together to provide the best performance and that removing some of them would cause performance degeneration or instability issues. From the results with , it is observed that all the template tokens (i.e., , , and “:”) contribute to the final performance. Removing one of them would cause a performance degradation. From the results without , the performance becomes less predictable, where adding back a template token (e.g., “:”) does not always bring performance improvements. Moreover, in some datasets, models without can still achieve relatively high performance. These results show that representations of other template tokens may also be seen as information anchors whose representations aggregate and serve information to the final prediction of LLMs, broadening the conclusions of wang2023label who claim that only answer tokens serve as information anchors.
Appendix E In-context Learning Templates
@@
Appendix F Representation-level Ablation Examples
We provide a one-shot demonstration of all ablation cases for the representation-level ablation experiments, as shown in Table 16. In these demonstrations, representations of specific token types are masked in the attention mechanism, where (m) denotes that all representations of the token are removed from the attention scope from the test example.
@@ Appendix F Representation-level Ablation Examples
| (s) Classify the news articles into the categories of World, Sports, Business, and Technology. |
| Article: (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) |
@@
| (s) Classify the news articles into the categories of World, Sports, Business, and Technology. |
| (m) (m) (m) (m) (m) (m) the (m) . (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) the (m) (m) (m) (m) (m) (m) (m) (m) , (m) (m) (m) (m) (m) (m) of (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) into (m) (m) the (m) (m) . |
| (m) (m) Technology |
@@
| (s) Classify the news articles into the categories of World, Sports, Business, and Technology. |
| (m) (m) (m) First class to (m) moon (m) London - British airline magnate Richard Branson announced a plan on Monday for (m) world’s first commercial space flights (m) saying “thousands” (m) fee-paying astronauts could be sent (m) orbit in (m) near future (m) (m) (m) (m) Technology (m) (m) |
@@
| (s) Classify the news articles into the categories of World, Sports, Business, and Technology. |
| (m) (m) (m) First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying ”thousands” of fee-paying astronauts could be sent into orbit in the near future. (m) (m) (m) Technology (m) (m) |
@@
| (s) Classify the news articles into the categories of World, Sports, Business, and Technology. |
| Article: First class to (m) moon (m) London - British airline magnate Richard Branson announced a plan on Monday for (m) world’s first commercial space flights (m) saying “thousands” (m) fee-paying astronauts could be sent (m) orbit in (m) near future (m) (m) Answer: Technology (m) (m) |
@@
| (s) Classify the news articles into the categories of World, Sports, Business, and Technology. |
| Article: (m) (m) (m) the (m) . (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) the (m) (m) (m) (m) (m) (m) (m) (m) , (m) (m) (m) (m) (m) (m) of (m) (m) (m) (m) (m) (m) (m) (m) (m) (m) into (m) (m) the (m) (m) . |
@@
Appendix G Experiment Results with More Large Language Models
Table 17 and 18 show the full results containing Llama 7B and Llama 13B models. Meanwhile, experimental results using the Llama 2 and Mistral models are shown in Table 19. The trends observed in these experiments are consistent with those involving the Llama models. These results further reinforce the findings of this paper, indicating that template tokens and stopword tokens are the most prone to serving as performance-critical tokens.
| Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 | 19.5 | ||
| cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 14.8 | |||
| stop | 36.7 | 82.9 | 32.0∗ | 52.4 | 58.8 | 56.2 | 33.7 | |||
| temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | 37.4 | |||
Zero-shot | 25.0 | 29.2 | 41.4 | 0.0 | 54.2 | 3.6 | 25.6 | ||
| cont | 32.4 | 57.9 | 42.5 | 12.5 | 55.5 | 46.1 | 15.6 | |||
| stop | 57.3 | 83.7 | 49.8 | 43.0 | 55.9 | 50.7 | 31.1 | |||
| temp | 70.8 | 90.2 | 58.4 | 66.2 | 66.3 | 73.5 | 45.3 | |||
Zero-shot | 59.0 | 18.0 | 37.0 | 0.0 | 0.0 | 0.0 | 19.0 | ||
| cont | 27.7 | 52.4 | 33.5 | 10.9 | 61.7 | 41.7 | 19.0 | |||
| stop | 72.2 | 73.5 | 46.8 | 50.7 | 58.6 | 30.6 | 36.4 | |||
| temp | 80.0 | 92.3 | 58.6 | 76.9 | 68.5 | 47.7 | 51.7 | |||
Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 | 54.6 | ||
| cont | 24.4 | 61.7 | 62.1 | 10.5 | 65.2 | 63.6 | 6.7 | |||
| stop | 72.9 | 92.7 | 66.7∗ | 69.1 | 69.6 | 63.0 | 17.7 | |||
| temp | 80.5 | 95.2 | 65.2 | 75.2 | 79.0 | 80.0 | 24.6 |
| Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | |||
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 | |||
| cont | 58.2 | 86.9 | 27.6 | 61.9 | 56.5 | 51.7 | 0.9 | ||||
| stop | 51.8 | 78.9 | 28.8 | 30.3 | 53.6 | 45.2 | 9.9 | ||||
| temp | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 23.8 | ||||
Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 | |||
| cont | 77.9 | 91.5 | 58.5 | 66.5 | 67.8 | 74.4 | 2.0 | ||||
| stop | 78.5 | 88.7 | 39.3 | 66.7 | 60.6 | 60.4 | 9.1 | ||||
| temp | 20.8 | 58.2 | 32.4 | 11.6 | 54.4 | 46.0 | 37.6 | ||||
Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 | |||
| cont | 81.4 | 93.1 | 58.9 | 75.7 | 69.6 | 45.1 | 0.2 | ||||
| stop | 79.8 | 85.8 | 64.4 | 73.6 | 64.5 | 40.2 | 2.4 | ||||
| temp | 27.8 | 52.4 | 33.5 | 10.9 | 63.1 | 45.6 | 31.5 | ||||
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 | |||
| cont | 82.3 | 95.4 | 64.9 | 76.1 | 80.4 | 82.0 | 1.5 | ||||
| stop | 84.8 | 94.9 | 62.1 | 77.3 | 70.5 | 74.4 | 4.3 | ||||
| temp | 24.4 | 61.7 | 60.6 | 10.5 | 67.7 | 68.5 | 32.7 |
| Models | Setting | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||
Zero-shot | 36.0 | 50.2 | 50.4 | 57.2 | 6.4 | 51.6 | 0.0 | ||
| cont | 1.9 | 0.9 | 61.0 | 50.6 | 12.9 | 48.7 | 53.2 | |||
| stop | 23.4 | 49.0 | 78.1 | 54.4 | 61.6 | 65.3 | 47.9 | |||
| temp | 31.3 | 81.1 | 82.6 | 55.2 | 65.5 | 63.9 | 55.4 | |||
Zero-shot | 53.3 | 56.2 | 90.8 | 49.0 | 7.6 | 70.0 | 46.4 | ||
| cont | 14.1 | 0.5 | 56.0 | 61.4 | 0.0 | 62.6 | 54.6 | |||
| stop | 9.6 | 47.2 | 76.8 | 65.2 | 65.3 | 66.5 | 56.8 | |||
| temp | 18.1 | 78.2 | 93.7 | 62.4 | 70.4 | 71.9 | 52.1 | |||
Zero-shot | 59.5 | 77.8 | 84.4 | 73.0 | 57.8 | 1.8 | 62.1 | ||
| cont | 10.0 | 43.3 | 52.0 | 66.6 | 10.1 | 64.3 | 60.7 | |||
| stop | 18.4 | 78.9 | 92.5 | 71.6 | 81.4 | 69.9 | 72.9 | |||
| temp | 19.7 | 81.7 | 95.9 | 63.9 | 83.3 | 77.9 | 72.5 | |||
Standard ICL | 70.7 | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 | ||
| cont | 3.8 | 82.4 | 85.5 | 54.3 | 64.2 | 59.6 | 55.7 | |||
| stop | 2.5 | 84.8 | 88.0 | 51.7 | 65.7 | 65.8 | 53.2 | |||
| temp | 32.8 | 0.9 | 61.0 | 50.6 | 12.9 | 48.5 | 53.6 | |||
Standard ICL | 73.6 | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 | ||
| cont | 1.3 | 79.0 | 94.1 | 62.7 | 72.4 | 72.1 | 53.6 | |||
| stop | 2.9 | 80.1 | 89.4 | 61.5 | 74.1 | 69.6 | 49.3 | |||
| temp | 33.5 | 0.5 | 56.0 | 61.4 | 0.0 | 68.2 | 54.3 | |||
Standard ICL | 80.2 | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 | ||
| cont | 0.6 | 81.8 | 96.2 | 64.4 | 83.4 | 78.9 | 72.9 | |||
| stop | 0.8 | 81.3 | 97.0 | 66.5 | 80.5 | 75.5 | 75.7 | |||
@@ Appendix G Experiment Results with More Large Language Models
Appendix H Experiment Results with Instruction-tuned Large Language Models
We present the representation-level ablation results of large language models after instruction tuning to confirm that our findings remain consistent. Specifically, we use Llama 2 7B Chat and Llama 2 13B Chat in our experiments. As shown in Table 20, the results align with the findings discussed in Section 5.2, with only one exception: the TREC dataset. In this dataset, the input data is structured in a question-answering format (e.g., Who/What/When did …). We hypothesize that, during the supervised fine-tuning process, tokens associated with these question formats may also serve as performance-critical tokens, although they are currently categorized as content tokens in our experiments. Overall, these supplemental results prove further evidence to our findings in the main paper.
| Models | Setting | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||
Zero-shot | - | - | - | - | - | - | - | ||
| cont | 27.9 | 0.7 | 52.6 | 52.2 | 7.9 | 24.8 | 28.9 | |||
| stop | 72.9 | 77.1 | 90.6 | 60.2 | 75.4 | 66.7 | 67.5 | |||
| temp | 75.6 | 80.1 | 92.6 | 62.6 | 76.6 | 69.5 | 72.1 | |||
Zero-shot | - | - | - | - | - | - | - | ||
| cont | 38.4 | 0.0 | 55.7 | 67.3 | 0.5 | 63.5 | 43.2 | |||
| stop | 67.1 | 78.5 | 87.6 | 66.9 | 71.4 | 67.2 | 31.1 | |||
| temp | 72.3 | 82.0 | 93.7 | 65.6 | 72.3 | 72.3 | 47.9 | |||
Standard ICL | - | - | - | - | - | - | - | ||
| cont | 76.1 | 80.7 | 93.1 | 62.9 | 76.8 | 70.7 | 72.5 | |||
| stop | 76.4 | 81.7 | 94.4 | 61.9 | 74.9 | 71.0 | 74.3 | |||
| temp | 31.4 | 0.7 | 52.6 | 52.2 | 7.9 | 24.1 | 51.1 | |||
Standard ICL | - | - | - | - | - | - | - | ||
| cont | 74.7 | 83.4 | 93.6 | 66.4 | 74.3 | 74.6 | 56.1 | |||
| stop | 69.8 | 78.9 | 94.1 | 57.7 | 72.8 | 70.1 | 45.4 | |||
@@ Appendix H Experiment Results with Instruction-tuned Large Language Models
Appendix I Stopword Tokens
For the results shown in the main paper, we used the stopword token (sarica2021stopwords) list shown in Table 11. This list only includes the stopword tokens from the task instruction, aiming to minimize their presence. We made this choice under the assumption that task-affecting information should be stored densely in a few tokens. Hence, the number of tokens whose representations affect the final task performance significantly should be small.
Nevertheless, one might be curious about the results if we used a more complete stopword list. In this case, we utilize a more comprehensive stopword token list of NLTK666https://gist.github.com/sebleier/554280 shown in Table 12 and conduct the representation-level ablation once more. The results are presented in Table 21. It can be observed that all the conclusions from Section 5.2 are still well established. A few results are different from Table 3 because we masked the representations of the “(/s)” token in this set of experiments. We claim that this masking does not impact the main findings of these experiments.
| Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||
Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 | ||
| + cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | |||
| + stop | 38.0 | 85.1 | 31.6 | 54.6 | 58.8 | 55.7 | |||
| + temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | |||
| Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | |||
| - temp | 42.1 | 87.2 | 25.9 | 56.3 | 58.3 | 57.4 | |||
| - cont | 57.1 | 88.4 | 27.1 | 62.6 | 56.8 | 52.4 | |||
| - stop | 61.6 | 90.7 | 24.8 | 62.2 | 56.7 | 51.9 | |||
Zero-shot | 25.0 | 29.2 | 41.4 | 0.0 | 54.2 | 3.6 | ||
| + cont | 32.4 | 57.9 | 42.5 | 12.5 | 55.5 | 46.1 | |||
| + stop | 59.9 | 85.9 | 51.7 | 28.9 | 56.0 | 52.7 | |||
| + temp | 70.8 | 90.2 | 58.4 | 66.2 | 66.3 | 73.5 | |||
| Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | |||
| - temp | 64.7 | 84.1 | 54.0 | 56.7 | 56.1 | 48.2 | |||
| - cont | 75.4 | 93.8 | 59.8 | 67.5 | 66.8 | 74.8 | |||
| - stop | 81.4 | 94.2 | 60.5 | 67.9 | 67.6 | 72.1 | |||
Zero-shot | 59.0 | 18.0 | 37.0 | 0.0 | 0.0 | 0.0 | ||
| + cont | 30.6 | 52.4 | 43.8 | 13.0 | 60.2 | 45.5 | |||
| + stop | 72.7 | 78.7 | 49.2 | 27.4 | 58.5 | 27.1 | |||
| + temp | 78.5 | 92.3 | 59.0 | 74.2 | 67.4 | 52.3 | |||
| Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | |||
| - temp | 71.7 | 80.1 | 56.2 | 8.7 | 56.5 | 29.3 | |||
| - cont | 79.3 | 93.4 | 60.1 | 74.1 | 68.4 | 47.6 | |||
| - stop | 79.2 | 94.1 | 59.3 | 73.8 | 68.9 | 44.6 | |||
Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 | ||
| + cont | 27.8 | 61.7 | 61.9 | 10.8 | 64.2 | 68.1 | |||
| + stop | 74.7 | 93.6 | 66.9 | 70.8 | 69.1 | 63.8 | |||
| + temp | 80.6 | 95.2 | 63.1 | 71.9 | 78.7 | 84.0 | |||
| Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | |||
| - temp | 79.5 | 93.8 | 58.5 | 62.8 | 68.0 | 68.0 | |||
| - cont | 82.7 | 95.9 | 62.9 | 74.1 | 79.6 | 83.1 | |||
| - stop | 84.4 | 96.1 | 61.8 | 72.8 | 79.4 | 82.1 |
Appendix J Analysis of the Next-line Tokens